I'm not sure I ma doing this right.
problem: (3/4)x*-1=1/2
I think i need to get the x by itself, so I would take the 3/4 and multiply it by 4/3, causing me to cancel it out. Leaving me with x-1=1/2. so then it's 1/1*1/2=1/2 so x=1/2? Is that right? I feel like I am missing something here.

Answer by rfer(16322)

3/4x*-1=1/2
3/4x=1/2*-1/1
3/4x=-1/2
x=-1/2*4/3
x=-4/6
x=-2/3